OpenStudy (4itc):

what fraction lies exactly halfway between 2/3 and 3/4 ? a) 3/5 b) 5/6 c) 7/12 d) 9/16 e) 17/24

OpenStudy (mathstudent55):

Find the average of 2/3 and 3/4. Add the two fractions, and divide by 2.

OpenStudy (4itc):

how do you find the average?

OpenStudy (4itc):

make the demonstrators the same?

OpenStudy (mathstudent55):

I explained above. Add the two fractions and divide by 2.

OpenStudy (4itc):

times one side by 4 and the other by 3= 8+9 /12 = 17/12= got crazy number then divide by 2 goth another crazy number then found the fraction is... e) 17/24

OpenStudy (mathstudent55):

 The sum of the fractions is 17/12. Now we need to divide 17/12 by 2.